# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v2

@@ E.1.2 Proof of Theorem 1
According to Proposition 6, any measure-zero event in system $F$ corresponds to a measure-zero even in system $\mathrm{GD}_{\eta}$. According to Proposition 8, the orbit of $\mathrm{GD}_{\eta}$ converges to $\left\{u^{\top}v=y\right\}$ if and only if the orbit of $F$ converges to $\{z=0\}$, and the former converges to $(\mathbf{0},\mathbf{0})$ if and only the latter converges to (y,0)(-y,0). According to Proposition 15, when $|\mu|<1$ and for almost all initializations $(z,w)$, the orbit converges to $\{z=0\}$ if $Q(z,w)<8$. Notice that $\mu=\eta y$ and due to the conjugacy (7),
$Q(z,w)<8$ $\Leftrightarrow\eta(\|u\|_{2}^{2}+\|v\|_{2}^{2})+\sqrt{\eta^{2}(\|u\|_{2}^{2}+\|v\|_{2}^{2})^{2}+16\eta y\cdot\eta(u^{\top}v-y)}<8$
$\Leftrightarrow(\|u\|_{2}^{2}+\|v\|_{2}^{2})+\sqrt{(\|u\|_{2}^{2}+\|v\|_{2}^{2})^{2}+16y\cdot(u^{\top}v-y)}<\frac{8}{\eta}.$
@@
We now show the sensitivity to initialization. Consider any open neighborhood $W\subset\mathbb{R}^{2d}$ such that $W\cap\partial\mathcal{D}^{\prime}_{\eta}\neq\varnothing$. Notice that the Jacobian of the map $T$ drops rank if and only if $\left\{u=\pm v\right\}$. Also, we have that
$(u,v)\in\partial\mathcal{D}_{\eta}^{\prime}$ $\Leftrightarrow\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2}+\sqrt{(\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2})^{2}-16y(u^{\top}v-y)}=\frac{8}{h}$
$\Rightarrow 4=\eta(\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2})+\eta^{2}y(u^{\top}v-y)$
$\Rightarrow(u^{\top}\ v^{\top})\begin{pmatrix}\eta I&\eta^{2}y/2\\ \eta^{2}y/2&\eta I\end{pmatrix}\begin{pmatrix}u\\ v\end{pmatrix}=4+\eta^{2}y^{2}.$
The last equation is a quadratic form, and the eigenvalues of the coefficient matrix are $\eta\pm\frac{1}{2}\eta^{2}|y|$, each with multiplicity $d$. Therefore, when $\eta|y|<1$, the quadratic form is positive definite and defines a smooth ellipsoid of dimension 2d12d-1. Notice that, $\left\{u=\pm v\right\}$ is the union of two linear subspace with dimension $d$. Therefore, since $2d-1\geq d$ and an ellipsoid is curved everywhere, there always exits a point $\bar{\theta}\in W\cap\partial\mathcal{D}^{\prime}_{\eta}\setminus\left\{u=\pm v\right\}$ and a neighborhood $\bar{W}$ of $\bar{\theta}$ such that $\bar{W}\subset W$ and $\bar{W}\cap\left\{u=\pm v\right\}=\varnothing$. The Jacobian of $T$ is full rank at all points in $\bar{W}$, so by constant rank theorem, $T(\bar{W})$ is an open set. Meanwhile, $T(\bar{\theta})\in T(\partial\mathcal{D}_{\eta}^{\prime})$. Hence, under the conjugacy (7), we have $T(\bar{W})\cap\left\{w=\mu z+4\right\}\neq\varnothing$. According to Proposition 20, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in T(\bar{W})$ such that $F^{N}(z^{\prime},w^{\prime})$ converges to $(0,w^{*})$ with any $w^{*}\in[2|\mu|,\infty)$ and $F^{N}(z^{\prime\prime},w^{\prime\prime})$ converges to $(-\mu,0)$, as $N$ tends to infinity. Therefore, there exists $\theta^{\prime},\theta^{\prime\prime}\in W$ such that $\mathrm{GD}_{\eta}^{N}(\theta^{\prime})$ converges to a global minimizer with squared norm in $2|y|,\infty)$ and $\mathrm{GD}_{\eta}^{N}(\theta^{\prime\prime})$ converges to $(\boldsymbol{0},\boldsymbol{0})$. Notice that
@@
Lastly, we present a basic property for the unregularized scalar factorization problem: the sharpness coincides with the squared norm at the set of global minimizers. This has been proved by Wang et al., (2022). Below we represent their results.
Proposition 21 (Wang et al.,, 2022, Theorem F.2).
@@ E.2.1 Preliminary results
We use $\mathcal{D}$ for $\mathcal{D}_{\eta}^{\prime\prime}$ for notation simplicity. First, we prove that $T(\partial\mathcal{D}_{\eta}^{\prime\prime})=\partial T(\mathcal{D_{\eta}^{\prime\prime}})$. Notice that, by Proposition 8, orbit under $\mathrm{GD}_{\eta}$ converges to $\mathcal{M}$ or to the saddle if and only if the corresponding orbit under $F$ converges to $T(\mathcal{M})$ or to $(-\mu,0)$, respectively. This implies that $T^{-1}(T(\mathcal{D}))=\mathcal{D}$ and $T(\mathcal{D}^{c})=T(\mathcal{D})^{c}$. Since $T$ is continuous, we have that $\partial T^{-1}(A)\subset T^{-1}(\partial A)$ for any set $A$. Hence, $\partial T^{-1}(T(\mathcal{D}))=\partial\mathcal{D}\subset T^{-1}(\partial T(\mathcal{D}))$. It follows that, $T(\partial\mathcal{D})\subset\partial T(\mathcal{D})$.
For the other direction, consider any point $y\in\partial T(\mathcal{D})$ and assume $y\notin T(\partial\mathcal{D})$ for contradiction. Then we have $T^{-1}(y)\cap\partial\mathcal{D}=\varnothing$. Note $T^{-1}(y)$ is a connected set by Proposition 7. Thus, there exists an open neighborhood $O$ of $T^{-1}(y)$ such that $O\subset\mathcal{D}^{o}$ or $O\subset(\mathcal{D}^{c})^{o}$. Hence, $T(O)\subset T(\mathcal{D})$ or $T(O)\subset T(\mathcal{D}^{c})=T(\mathcal{D})^{c}$, which contradicts the fact that $y\in\partial T(\mathcal{D})$. Hence $y\in T(\partial\mathcal{D})$ and $\partial T(\mathcal{D})\subset T(\partial\mathcal{D})$. It follows that $\partial T(\mathcal{D})=T(\partial\mathcal{D})$.
Next, we prove that, $F(\partial T(\mathcal{D}))=\partial T(\mathcal{D})$. Now let $A=T(\mathcal{D})$ for simplicity. Note that $A$ satisfies $F(A)\subset A$, $F^{-1}(A)\subset A$, since $A=T(\mathcal{D})$ contains all initializations that converge to $T(\mathcal{M})$ and to $(-\mu,0)$. Consider any point $x\in F^{-1}(\partial A)$. Then a small enough neighborhood $O$ of $x$ is mapped to a neighborhood of $F(x)$, which contains a point $y\in A$ and a point $z\in A^{c}$. Since $F^{-1}(A)\subset A$, $F^{-1}(y)\in O\cap A$. Since $F(A)\subset A$, $F^{-1}(z)\in O\cap A^{c}$. Therefore, $x\in\partial A$ and $F^{-1}(\partial A)\subset\partial A$. Meanwhile, since $F$ is surjective by Proposition 13, we have $F\circ F^{-1}(\partial A)=\partial A$. It follows that
@@
For the other direction, note that for any $y\in\partial A$, since $y\in\mathrm{cl}(A)$ and $F$ is continuous, we have that $F(y)\in\mathrm{cl}(F(A))\subset\mathrm{cl}(A)$. Consider the set $\left\{Q(z,w)<8-4\nu\right\}$. By Proposition 22, $\left\{Q(z,w)<8-4\nu\right\}$ is an open neighborhood of $T(\mathcal{M})\cup\{(\mu,0)\}$ and all orbits in this neighborhood converge to $T(\mathcal{M})\cup\{(\mu,0)\}$. Now consider any $x\in A$. Since $F^{N}(x)$ converges to $T(\mathcal{M})$ or to $(\mu,0)$, $F^{N}(x)$ must visit $\Gamma$ at some $N$. Since $\Gamma$ is open and $F^{N}$ is continuous, $x$ is an interior point of $A$. Thus, $A$ is an open set. Therefore, $y\in\mathrm{cl}(A)\setminus A^{o}=\mathrm{cl}(A)\setminus A$. Since $F^{-1}(A)\subset A$, $F(y)\notin A$. Taken together, we have that $F(y)\in\mathrm{cl}(A)\setminus A=\partial A$ and $F(\partial A)\subset\partial A$. Therefore, $F(\partial A)=\partial A$.
By proposition 22, we have that $A$ contains $\left\{Q(z,w)<8-4\nu\right\}$. Therefore, its boundary, $\partial A$, must lie in $\left\{Q\geq 6-4\nu\right\}$. Since $F(\partial A)=\partial A$, by Proposition 13, we have that
@@
Assume $0\leq\nu<\min\{\frac{1}{2},1-|\mu|\}$. Consider arbitrary point $m_{0}=(z,w)\in\Omega$. When $\mu\geq 0$, $\lim_{N\to\infty}G_{0}^{N}(m_{0})=(-2+\nu,4-2(\nu+\mu))$. When $\mu<0$, $\lim_{N\to\infty}G_{2}^{N}(m_{0})=(2-\nu,4+2(\nu+\mu))$.
@@
Assume $0\leq\nu<\min\{\frac{1}{2},1-|\mu|\}$. Consider $\xi_{1}=(-2+\nu,4-2(\nu+\mu))$ and $\xi_{2}=(2-\nu,4+2(\nu+\nu))$. For $i=1,2$, we have that $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})$ has infinitely many points and $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})\subset T(\partial\mathcal{D}_{\eta}^{\prime\prime})$. When $y\geq 0$, for any open set $O$ such that $O\cap(\cup_{N=0}^{\infty}F^{-N}(\xi_{1}))\neq\varnothing$, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in O$ such that, for any $(u^{\prime},v^{\prime}),(u^{\prime\prime},v^{\prime\prime})$ that satisfy $T(u^{\prime},v^{\prime})=(z^{\prime},w^{\prime})$ and $T(u^{\prime\prime},v^{\prime\prime})=(z^{\prime\prime},w^{\prime\prime})$, we have $\mathrm{GD}_{\eta}^{N}(u^{\prime},v^{\prime})$ converges to $p^{+}(u^{\prime},v^{\prime})$ and $\mathrm{GD}_{\eta}^{N}(u^{\prime\prime},v^{\prime\prime})$ converges to $p^{-}(u^{\prime\prime},v^{\prime\prime})$. When $y<0$, the same result holds for any open set $O$ such that $O\cap(\cup_{N=0}^{\infty}F^{-N}(\xi_{2}))\neq\varnothing$.
@@
Next, we show that for any open set $O$ such that $O\cap\cup_{N=0}^{\infty}F^{-N}(\xi_{1})\neq\varnothing$, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in O$ satisfying the claimed properties. Note, by (25), it suffices to prove this for $O$ that satisfies $O\ni\xi_{1}$. When $y\leq\lambda$, $L$ has a unique minimizer $(\boldsymbol{0},\boldsymbol{0})$ and the result obviously holds from Proposition 25. Now consider $y>\lambda$. Let $(u_{k},v_{k})_{k\geq 0}$ be a gradient descent orbit that converges to a global minimizer, and $(z_{k},w_{k})=T(u_{k},v_{k})$. As we shown in the proof of Proposition 22, the converged minimizer is the minimal distance solution if
$\left\{j\geq 0\colon 1-z_{j}-\nu<0\right\}$
is an even number; and the converged minimizer is the maximal distance solution if the above is an odd number. Now let $m_{*}=(-\nu,2(\mu-\nu))$, where $\{m^{*}\}=T(\mathcal{M})$. Note, the boundary $\{w=2(z+\mu)\}$ is invariant under $F$ by Lemma 11, where the $w$-update under $F$ is given by
$w^{\prime}=d(w)=w(\frac{w}{2}-1-\mu+\nu)^{2}.$
Solving $d(w)=2(\mu-\nu)$ yields two solutions:
$w_{\pm}=2+\mu-\nu\pm\sqrt{-4+(-2-\mu+\nu)^{2}}.$
Note these roots are reals since $\mu-\nu>0$. Note on the line $\{w=2(z+\mu)\}$, $1-\nu=z$ yields $w=2(1-\nu+\mu)$. Let $\alpha=\mu-\nu$. We have that
$2(1+\alpha)<2+\alpha+\sqrt{(2+\alpha)^{2}-4}$
[…]
which is true by assumption. Similarly, we have
$2(1+\alpha)>2+\alpha-\sqrt{(2+\alpha)^{2}-4}$
[…]
which is also true as $\alpha>0$. Therefore, there exists $m_{*}^{+},m_{*}^{-}\in\Omega$ such that, $F(m_{*}^{\pm})=m^{*}$ and $m_{*}^{+}\in\left\{z>1-\nu\right\}$ and $m_{*}^{-}\in\left\{z<1-\nu\right\}$. By Proposition 25, $G_{0}^{N}(m_{*}^{\pm})$ converges to $(2-\nu,4+2(\nu+\mu))$ as $N\to\infty$. This means that we can find $(z^{\prime},w^{\prime})$ and $(z^{\prime\prime},w^{\prime\prime})\in O$ such that $(z^{\prime},w^{\prime})=G_{0}^{N^{\prime}}(m_{*}^{+})$ and $(z^{\prime\prime},w^{\prime\prime})=G_{0}^{N^{\prime\prime}}(m_{*}^{-})$ for some $N^{\prime},N^{\prime\prime}>0$. Notice that, for $N\geq 1$, $G_{0}^{N}(m_{*}^{\pm})\in\left\{z<1-\nu\right\}$ since the image set of $G_{0}$ is $\mathrm{cl}(A_{0})$. Therefore, the $F$-orbit initialized at $(z^{\prime},w^{\prime})$ enters the region $\{z>1-\nu\}$ exactly once, while the $F$-orbit initialized at $(z^{\prime\prime},w^{\prime\prime})$ never enters $\{z>1-\nu\}$. Therefore, the corresponding $uv$-orbit converges to the maximal and minimal distance solutions, respectively. This completes the proof. ∎

### v3

@@ E.1.2 Proof of Theorem 1 and Theorem 2
According to Proposition 7, any measure-zero event in system $F$ corresponds to a measure-zero event in system $\mathrm{GD}_{\eta}$. According to Proposition 9, the orbit of $\mathrm{GD}_{\eta}$ converges to $\left\{u^{\top}v=y\right\}$ if and only if the orbit of $F$ converges to $\{z=0\}$, and the former converges to $(\boldsymbol{0},\boldsymbol{0})$ if and only the latter converges to (y,0)(-y,0). According to Proposition 16, when $|\mu|<1$ and for almost all initializations $(z,w)$, the orbit converges to $\{z=0\}$ if $Q(z,w)<8$. Notice that $\mu=\eta y$ and due to the conjugacy (8),
$Q(z,w)<8$ $\Leftrightarrow\eta(\|u\|_{2}^{2}+\|v\|_{2}^{2})+\sqrt{\eta^{2}(\|u\|_{2}^{2}+\|v\|_{2}^{2})^{2}-16\eta y\cdot\eta(u^{\top}v-y)}<8$
$\Leftrightarrow\|u\|_{2}^{2}+\|v\|_{2}^{2}+\sqrt{(\|u\|_{2}^{2}+\|v\|_{2}^{2})^{2}-16y\cdot(u^{\top}v-y)}<\frac{8}{\eta}.$
@@
We now show the sensitivity to initialization. Consider any open neighborhood $W\subset\mathbb{R}^{2d}$ such that $W\cap\partial\mathcal{D}^{\prime}_{\eta}\neq\varnothing$. Notice that the Jacobian of the map $T$ drops rank if and only if $\left\{u=\pm v\right\}$. Also, as shown in (23), $\partial\mathcal{D}_{\eta}^{\prime}$ defines a smooth ellipsoid of dimension 2d12d-1. Notice that, $\left\{u=\pm v\right\}$ is the union of two linear subspace with dimension $d$. Therefore, since $2d-1\geq d$ and an ellipsoid is curved everywhere, there always exists a point $\bar{\theta}\in W\cap\partial\mathcal{D}^{\prime}_{\eta}\setminus\left\{u=\pm v\right\}$ and a neighborhood $\bar{W}$ of $\bar{\theta}$ such that $\bar{W}\subset W$ and $\bar{W}\cap\left\{u=\pm v\right\}=\varnothing$. The Jacobian of $T$ is full rank at all points in $\bar{W}$, so by constant rank theorem, $T(\bar{W})$ is an open set. Meanwhile, $T(\bar{\theta})\in T(\partial\mathcal{D}_{\eta}^{\prime})$. Hence, under the conjugacy (8), we have $T(\bar{W})\cap\left\{w=\mu z+4\right\}\neq\varnothing$. According to Proposition 21, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in T(\bar{W})$ such that $F^{N}(z^{\prime},w^{\prime})$ converges to $(0,w^{*})$ with any $w^{*}\in[2|\mu|,\infty)$ and $F^{N}(z^{\prime\prime},w^{\prime\prime})$ converges to $(-\mu,0)$, as $N$ tends to infinity. Therefore, there exists $\theta^{\prime},\theta^{\prime\prime}\in W$ such that $\mathrm{GD}_{\eta}^{N}(\theta^{\prime})$ converges to a global minimizer with squared norm in $2|y|,\infty)$ and $\mathrm{GD}_{\eta}^{N}(\theta^{\prime\prime})$ converges to $(\boldsymbol{0},\boldsymbol{0})$. Notice that
@@
Next, we present a basic property for the unregularized scalar factorization problem: the sharpness coincides with the squared norm at the set of global minimizers. This has been proved by Wang et al. (2022). Below we restate their results.
Proposition 22 (Wang et al., 2022, Theorem F.2).
@@
Proof of Theorem 2.
Assume $E=(e,e^{\prime})\subset[\gamma_{\min},2/\eta]$. Let $E^{\prime}=(e\eta,e^{\prime}\eta)$. We first prove that, in the $F$-system, there exists a positive-measure set of initializations that converge to $\left\{(0,w)\colon w\in E^{\prime}\right\}$. Notice that
$|\frac{z_{k+1}}{z_{k}}|=|z_{k}^{2}+\mu z_{k}-w_{k}+1|\leq 1$ (24)
is equivalent to
$w_{k}\geq z_{k}^{2}+\mu z_{k},\ \text{and}\ w_{k}\leq z_{k}^{2}+\mu z_{k}+2.$
Notice that (i) the convex parabola $w=z^{2}+\mu z+2$ intersects the $w$-axis at $(0,2)$, (ii) the level set $\left\{Q(z,w)=c\right\}$ are straight lines intersecting the $w$-axis at $(0,c/2)$. Therefore, there exists $\delta\in(0,1),C\in(4|\mu|,4)$ such that all points in
$T=\left\{(z,w)\in\Omega\colon|z|<\delta,Q(z,w)<C\right\}$
satisfies $|z_{k+1}/z_{k}|<\gamma$ for some constant $\gamma\in(0,1)$. Notice that, this holds for all sufficiently small $\delta$ and thus we fix $\gamma$. For $(z_{k},w_{k})\in T$ we have that $|z_{k+1}|<\delta$. Meanwhile, since $Q(z_{k+1},w_{k+1})\leq Q(z_{k},w_{k})$ by Lemma 13, we have that $(z_{k+1},w_{k+1})\in T$. It follows that if the initialization $(z_{0},w_{0})$ is in $T$, the entire trajectory remains within $T$. Hence, we have that $|z_{k}|\leq|z_{0}|\gamma^{k}$ for all $k\geq 0$, and that
$|w_{k+1}-w_{k}|=|z_{k}^{2}(w_{k}-4)-4z_{k}\mu|\leq|z_{k}|^{2}|w_{k}-4|+4|\mu||z_{k}|\leq C_{1}|z_{k}|$
where $C_{1}$ is independent of $\delta$. We have that,
$|w_{\infty}-w_{0}|\leq C_{1}\sum_{k=0}^{\infty}|z_{k}|\leq\frac{C_{1}\delta}{1-\gamma}.$
Hence, $|w_{\infty}-w_{0}|\to 0$ as $\delta\to 0$. Then, there must exist a sufficiently small sub-interval $E^{\prime\prime}\subset E^{\prime}$ and a sufficiently small $\delta$ such that, for all initializations in
$A=\left\{(z,w)\in\Omega\colon|z|<\delta,|w|\in E^{\prime\prime}\right\},$
which is a positive-measure set, the trajectory converges to $\left\{(0,w)\colon w\in E^{\prime}\right\}$.
Consider any open neighborhood $W\subset\mathbb{R}^{2d}$ such that $W\cap\partial\mathcal{D}^{\prime}_{\eta}\neq\varnothing$. As shown in the above proof of Theorem 1, there always exits a point $\bar{\theta}\in W\cap\partial\mathcal{D}^{\prime}_{\eta}\setminus\left\{u=\pm v\right\}$ and a neighborhood $\bar{W}$ of $\bar{\theta}$ such that $\bar{W}\subset W$, $T$ has full rank on $\bar{W}$, and $T(\bar{W})$ is an open set satisfying $T(\bar{W})\cap\left\{w=\mu z+4\right\}\neq\varnothing$. According to Proposition 21, there exists $(z^{\prime},w^{\prime})\in T(\bar{W})$ such that $F^{N}(z^{\prime},w^{\prime})\in A$. By continuity of $F^{N}$, there exists a neighbor $A^{\prime}$ of $(z^{\prime},w^{\prime})$ satisfying $A^{\prime}\subset T(\bar{W})$ and $F^{N}(A^{\prime})\subset A$. Since $T|_{\bar{W}}\colon\bar{W}\to T(\bar{W})$ is continuous, $T|_{\bar{W}}^{-1}(A^{\prime})$ is open in $\bar{W}$ and has positive measure. By the conjugacy (8) and what we have shown, for all points in $T|_{\bar{W}}^{-1}(A^{\prime})$, the final norm lies in $E$. This completes the proof. ∎
Lastly, we discuss the technical novelties in the proof of Theorem 1, in comparison with the techniques of Wang et al. (2022). While their analysis is also based on a careful study of the gradient descent system, it is restricted to a specific subset of the parameter space. In comparison, our approach is built on a global analysis of the system. The key ingredients are: (i) we identified an equivariance symmetry in $\mathrm{GD}_{\eta}$ and the corresponding quotient dynamical system defined by the map $F$ (Proposition 9); (ii) we analyzed all the branches of the inverse map $F^{-1}$ (which is a multi-valued map) and carefully studied the inverse dynamics defined by $F^{-1}$ (Proposition 14 and Proposition 21); (iii) we used symbolic dynamics to show that $\mathrm{GD}_{\eta}$, when restricted to the convergence boundary, is semi-conjugate to a Devaney chaotic system (Proposition 18) and that $\mathrm{GD}_{\eta}$ admits periodic orbits with any period (Proposition 19).
@@ E.2.1 Preliminary results
We use $\mathcal{D}$ for $\mathcal{D}_{\eta}^{\prime\prime}$ for notation simplicity. First, we prove that $T(\partial\mathcal{D})=\partial T(\mathcal{D})$. Notice that, by Proposition 9, orbit under $\mathrm{GD}_{\eta}$ converges to $\mathcal{M}$ or to the saddle if and only if the corresponding orbit under $F$ converges to $T(\mathcal{M})$ or to $(-\mu,0)$, respectively. This implies that $T^{-1}(T(\mathcal{D}))=\mathcal{D}$ and $T(\mathcal{D}^{c})=T(\mathcal{D})^{c}$. Since $T$ is continuous, we have that $\partial T^{-1}(T(\mathcal{D}))=\partial\mathcal{D}\subset T^{-1}(\partial T(\mathcal{D}))$. Hence, $T(\partial\mathcal{D})\subset\partial T(\mathcal{D})$.
For the other direction, consider any point $y\notin T(\partial\mathcal{D})$. Then we have $T^{-1}(y)\cap\partial\mathcal{D}=\varnothing$. Note $T^{-1}(y)$ is connected and compact by Proposition 8. Thus, there exists an open neighborhood $O$ of $T^{-1}(y)$ such that $O\subset\mathcal{D}$ or $O\subset\mathcal{D}^{c}$. Hence, $T(O)\subset T(\mathcal{D})$ or $T(O)\subset T(\mathcal{D}^{c})=T(\mathcal{D})^{c}$. We claim that for any $y^{\prime}$ that is sufficiently close to $y$, $y^{\prime}\in T(O)$. Notice that, if $y\notin\partial\Omega$, we have that $T^{-1}(y)\cap\left\{u=\pm v\right\}=\varnothing$. Thus, $JT$ is non-singular at $T^{-1}(y)$ and, by constant rank theorem, $T(O)$ contains a neighborhood of $y$, which gives the claim. If $y\in\partial\Omega$, without loss of generality, assume that $y=(\frac{w_{0}}{2}-\mu,w_{0})$ for some $w_{0}\geq 0$. Then $F^{-1}(y)=\left\{u=v,\|u\|^{2}=w_{0}/2\right\}$. Consider any $(u^{\prime},v^{\prime})\in T^{-1}(y^{\prime})$. As shown in the proof of Proposition 8, when $\|y^{\prime}-y\|$ tends to zero, $\|u^{\prime}-v^{\prime}\|^{2}$ must tend to 0 and $\|u^{\prime}+v^{\prime}\|^{2}$ to $2w_{0}$. This implies that $(u^{\prime},v^{\prime})$ tends to $F^{-1}(y)$. Thus, if $y^{\prime}$ is sufficiently close to $y$, we have $T^{-1}(y)\subset O$ and thus $y^{\prime}\in T(O)$. Now, given that $T(O)\subset T(\mathcal{D})$ or $T(O)\subset T(\mathcal{D})^{c}$, we have $y\notin\partial T(\mathcal{D})$. Hence, $\partial T(\mathcal{D})\subset T(\partial\mathcal{D})$ and $\partial T(\mathcal{D})=T(\partial\mathcal{D})$.
Next, we prove that $F(\partial T(\mathcal{D}))=\partial T(\mathcal{D})$. Let $A=T(\mathcal{D})$. Notice that, by Proposition 23, $\left\{Q(z,w)<8-4\nu\right\}$ is an attracting neighborhood of $T(\mathcal{M})\cup\{(\mu,0)\}$. Thus, the corresponding basin of attraction, $A$, is an open set due to the continuity of $F$. It follows that $\partial A\subset A^{c}\subset\left\{Q\geq 8-4\nu\right\}$. Now consider any point $x\in F^{-1}(\partial A)$. We claim that a small enough neighborhood $O$ of $x$ is mapped to a neighborhood of $F(x)$. Note the claim holds trivially if $x$ is a regular point of $F$. Assume that $\det JF(x)=0$ and recall that the critical points of $F$ is given in (14). Since $F(x)\in\partial A\subset A^{c}$ and $F^{-1}(A^{c})\subset A^{c}$, we have that $x\in A^{c}\subset\left\{Q\geq 8-4\nu\right\}$. Hence, as we shown in the proof of Proposition 14, $x\in\partial A_{0}$ or $x\in\partial A_{2}$, and the claim holds according to Proposition 14. Now, since $F(x)\in\partial A$, $F(O)$ contains a point $y\in A$ and a point $z\in A^{c}$. By the previous claim, $F^{-1}(y)\cap O\neq\varnothing$ and $F^{-1}(z)\cap O\neq\varnothing$. As $F^{-1}(A)\subset A$ and $F^{-1}(A^{c})\subset A^{c}$, $x\in\partial A$ and $F^{-1}(\partial A)\subset\partial A$. Since $F$ is surjective by Proposition 14, we have $F\circ F^{-1}(\partial A)=\partial A$. It follows that
@@
For the other direction, note that for any $y\in\partial A$, since $y\in\mathrm{cl}(A)$ and $F$ is continuous, we have that $F(y)\in\mathrm{cl}(F(A))\subset\mathrm{cl}(A)$. Therefore, $y\in\mathrm{cl}(A)\setminus A^{o}=\mathrm{cl}(A)\setminus A$. Since $F^{-1}(A)\subset A$, $F(y)\notin A$. Then we have that $F(y)\in\mathrm{cl}(A)\setminus A=\partial A$ and $F(\partial A)\subset\partial A$. Therefore, $F(\partial A)=\partial A$.
Since $F(\partial A)=\partial A$ and $\partial A\subset\left\{Q\geq 8-4\nu\right\}$, Proposition 14 gives that
@@
Assume $0\leq\nu<\min\{\frac{1}{2},1-|\mu|\}$. Consider arbitrary point $m_{0}=(z,w)\in\Omega$. When $\mu\geq 0$, $\lim_{N\to\infty}G_{0}^{N}(m_{0})=(-2+\nu,4-2(\nu+\mu))$. When $\mu<0$, $\lim_{N\to\infty}G_{2}^{N}(m_{0})=(2-\nu,4+2(\mu-\nu))$.
@@
Assume $0\leq\nu<\min\{\frac{1}{2},1-|\mu|\}$. Consider $\xi_{1}=(-2+\nu,4-2(\nu+\mu))$ and $\xi_{2}=(2-\nu,4+2(\mu-\nu))$. For $i=1,2$, we have that $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})$ has infinitely many points and $\cup_{N=0}^{\infty}F^{-N}(\xi_{i})\subset T(\partial\mathcal{D}_{\eta}^{\prime\prime})$. When $y\geq 0$, for any open set $O$ such that $O\cap(\cup_{N=0}^{\infty}F^{-N}(\xi_{1}))\neq\varnothing$, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in O$ such that, for any $(u^{\prime},v^{\prime}),(u^{\prime\prime},v^{\prime\prime})$ that satisfy $T(u^{\prime},v^{\prime})=(z^{\prime},w^{\prime})$ and $T(u^{\prime\prime},v^{\prime\prime})=(z^{\prime\prime},w^{\prime\prime})$, we have $\mathrm{GD}_{\eta}^{N}(u^{\prime},v^{\prime})$ converges to $p^{+}(u^{\prime},v^{\prime})$ and $\mathrm{GD}_{\eta}^{N}(u^{\prime\prime},v^{\prime\prime})$ converges to $p^{-}(u^{\prime\prime},v^{\prime\prime})$. When $y<0$, the same result holds for any open set $O$ such that $O\cap(\cup_{N=0}^{\infty}F^{-N}(\xi_{2}))\neq\varnothing$.
@@
Next, we show that for any open set $O$ such that $O\cap\cup_{N=0}^{\infty}F^{-N}(\xi_{1})\neq\varnothing$, there exists $(z^{\prime},w^{\prime}),(z^{\prime\prime},w^{\prime\prime})\in O$ satisfying the claimed properties. When $y\leq\lambda$, $L$ has a unique minimizer $(\boldsymbol{0},\boldsymbol{0})$ and the result holds according to Proposition 26. Now consider $y>\lambda$. Let $(u_{k},v_{k})_{k\geq 0}$ be a gradient descent orbit that converges to a global minimizer, and $(z_{k},w_{k})=T(u_{k},v_{k})$. As we shown in the proof of Proposition 23, the converged minimizer is the minimal distance solution if
$\zeta(z_{0},w_{0})=\#\left\{j\geq 0\colon 1-z_{j}-\nu<0\right\}$
is an even number; and the converged minimizer is the maximal distance solution if the above is an odd number.
Let $m_{*}=(-\nu,2(\mu-\nu))$, so that $\{m_{*}\}=T(\mathcal{M})$. Assume that $O\ni\left\{G_{i_{1}}\circ\cdots\circ G_{i_{k}}(\xi_{1})\right\}$ for some fixed $i_{1},\cdots,i_{k}$. By the continuity of $G_{i_{1}}\circ\cdots\circ G_{i_{k}}(\xi_{1})$, there exists a neighborhood $\tilde{O}\ni\xi_{1}$ such that $G_{i_{1}}\circ\cdots\circ G_{i_{k}}(\xi_{1})(\tilde{O})\subset O$. By Proposition 26, there exists $N^{\prime}>0$ such that $G_{0}^{N^{\prime}}(\Omega)\subset\tilde{O}$. Then we have that
$(z^{\prime},w^{\prime})\triangleq G_{i_{1}}\circ\cdots\circ G_{i_{k}}\circ G_{0}^{N^{\prime}}(m_{*})\in O$
and
$(z^{\prime\prime},w^{\prime\prime})\triangleq G_{i_{1}}\circ\cdots\circ G_{i_{k}}\circ G_{0}^{N^{\prime}}\circ G_{2}(m_{*})\in O.$
By construction, $F^{k+N^{\prime}}(z^{\prime},w^{\prime})=m_{*}$ and $F^{k+N^{\prime}+1}(z^{\prime\prime},w^{\prime\prime})=m_{*}$. Then it suffices to that one of $\zeta(z^{\prime},w^{\prime})$ and $\zeta(z^{\prime\prime},w^{\prime\prime})$ is odd and the other is even. To see this, notice that $\left\{w=-2(z+\mu)\right\}$ is forward-invariant. Then, as $m^{*}\notin\left\{w=-2(z+\mu)\right\}$, the orbit starting from $(z^{\prime},w^{\prime})$ and from $(z^{\prime\prime},w^{\prime\prime})$ can not visit $\left\{w=-2(z+\mu)\right\}$. Therefore, by noticing that $G_{0}(\Omega)\subset\left\{z<1-\nu\right\}$, $G_{1}(\Omega\setminus\left\{w=-2(z+\mu)\right\})\subset\left\{z<1-\nu\right\}$, $G_{2}(\Omega\setminus\left\{w=-2(z+\mu)\right\})\subset\left\{z>1-\nu\right\}$, and $m_{*}\in\left\{z<1-\nu\right\}$, we have that
$\zeta(z^{\prime},w^{\prime})=\#\left\{1\leq j\leq k\colon i_{j}=2\right\},\ \zeta(z^{\prime\prime},w^{\prime\prime})=\#\left\{1\leq j\leq k\colon i_{j}=2\right\}+1.$
This completes the proof. ∎